OpenStudy (anonymous):

If 2 cards are drawn from a deck of 52 cards, what is the expected number of spades? A. 0.75 B. 0.25 C. 0.50 D. 0.47

OpenStudy (anonymous):

B

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

probability you get a spade is in fact \(\frac{1}{4}\) but you are asked for expected value, not the probability

OpenStudy (anonymous):

\[\frac{1}{4}\times 2=\frac{1}{2}=.5\]